radomes.openskiesproject.org
cinthiagonzales cinthiagonzales
  • 15-02-2017
  • Mathematics
Answered

an experiment consists of flipping a fair coin 4 times. What is the probability of obtaining at least one head?

Verified Answer :

LammettHash
LammettHash LammettHash
  • 16-02-2017

\(\mathbb P(X\ge1)=1-\mathbb P(X=0)\)

The probability of getting 0 heads in 4 tosses (or equivalently, 4 tails) is \(\dfrac1{2^4}=\dfrac1{16}\).

So the desired probability is

\(1-\dfrac1{16}=\dfrac{15}{16}\)

Answer Link

Related Questions

Which best identifies a runon sentenceA A lioncan run faster than a greyhound or a foxB Humpback whales sing songs every yearC some tortoises are huge you can r
solve the equation for y y32 4y12
How did humanism influence the works of Renaissance artists and authors
how did the English overcome the Spanish armada
7 Which one of the following statements is most accurate regarding the Mayflower Compact A It created the Massachusetts Bay Colony B It founded the Plymouth col
What is the standard form for 200000800007006
Which of these people is most likely to experience stress Sheila has two jobs tutoring Spanish and catering for weddings If she spent less time tutoring she cou
In Glaspell's The Inheritors, which quotation best supports the feminist theme that women often feel a desire to control their own fate? A. MADELINE: He got
One reason President George Washington called for a policy of neutrality in the 1790s was to 1 improve his popularity among Federalist voters 2 avoid United Sta
gravity will have a greater effect on an object with 2 points less force more mass less weight more charge